# Product-oriented Product–Process–Resource Asset Network and its Representation in AutomationML for Asset Administration Shell

### v1

@@ IV Representation of PoPAN in AutomationML
The AutomationML representations aims at capturing the PoPAN network as a graph. First, we specified the InterfaceClassLib with the Interface Class EdgeVertexInterface. This interface allows to connect edges and vertices via AutomationML Internal Links. The Interface Class has an attribute Direction, which can have values In for incoming edge, Out for outgoing edge, and InOut for undirected edge. Since PoPAN distinguishes assembly and disassembly directions, the EdgeVertexInterface Class has also attribute Arrow, which can have values OpenArrow for the assembly direction, FullArrow for the disassembly direction, or None for none arrow.
@@
In the third step, we have defined exemplary/generic System Unit Class, which are depicted in Fig. 2. The figure shows four Internal Elements (IE) from the System Unit Class, in compliance to their Roles from the Role Class Library. The Internal Elements are as a matter of fact instances of the aforementioned Role Classes, accompanied by EdgeVertexInterface to enable the PoPAN representation.
@@
The IE Product type represents physical products within the whole product. Each IE Product is uniquely identified by an Id and has a Type including ElementaryProduct, SubProduct, FastenerProduct, InitialProduct, or LastProduct. Additionally, IE Product has assigned Roles Product and Submodel (to corespond with AAS). The IE Process type represents processes performed within the assembly/disassembly operations. IE Process is uniquely identified by an Id, has assigned Role Process and has a Type. The IE Resource type represents resources needed to perform processes, such as tools and stations. Similar to IE Process, each IE Resource has a unique identifier Id, has assigned Role Process and has a Type. The IE Edge type represents connections between PPR components within the whole product. IE Edge is uniquely identified by Id, the Role Edge and has a Type that can have values Physical to represent the physical continuity of products, Structural for the structural continuity of products, LinkingP2P for linking a product to a process, SequenceProcess to represent dependencies/sequences of processes, and LinkingP2R for linking a process to a resource. All four IE Product types The EdgeVertexInterface is defined to ensure connection to other IEs within the PoPAN.
@@
In addition, the Internal Element Product supports the role Submodel from the Asset Administration Shell Role Class Library [^3] Online: https://www.automationml.org/wp-content/uploads/2022/04/Asset-Administration-Shell-Representation-V1_0_0.zip. In this place we should note that we are missing a “SubModelElement/Entity” in this AutomationML library, which would be useful for transforming the description into the AASX format.
@@
The example of a generic product structure according to the PoPAN model from the previous section is depicted in Fig. 3. If we need to assign more than one edge to a vertice, we have to increase the number of EdgeVertexInterfaces. In Fig. 4, a network of PoPAN elements represented in AutomationML is shown.
@@
AutomationML supports annotating objects relevant for the Asset Administration Shell (AAS). AAS is becoming an important industrial artifact to be exchanged in the entire supply chain. For creating the AAS, we are using the aforementioned tool AASX Package Explorer, which can import an AutomationML file and transform it into the AASX format, which can accompany the product during its entire life-cycle.
@@
This comprehensive product description, when used in conjunction with the AAS, can accompany the product throughout its entire lifecycle. It can assist in scenarios where defective parts need replacement, while also facilitating the tracking of new or missing products (e.g., a missing screw) to streamline the recycling process.
@@ V Electric Vehicle Battery Remanufacturing Use-Case
In our laboratory conditions, utilizing the Industry 4.0 Testbed [^4] Online: https://www.ciirc.cvut.cz/teams-labs/testbed/ at CTU in Prague – CIIRC, we employ KUKA robots to manipulate the real EV battery according to predefined parameters (see Fig. 5). The primary goal of this use-case is to disassemble the entire EV battery by a flexible robotic system. This approach offers the potential use of modules from the EV battery for secondary applications. Additionally, the secondary goal involves replacing broken modules, which necessitates a combination of assembling and disassembling operations. We were seeking a formalism that allows to accommodate all possible operations in our use-case, and PoPAN fits our requirements perfectly for all considered applications. The visualization of PoPAN for a simplified description of the EV battery is depicted in Fig. 6. It illustrates all relevant data for creating assembly/disassembly operations. As outlined in Section III, this visualization contains product, processes, and resources interconnected by different types of edges, with two types of arrows indicating assembly/disassembly orientations. Following the rules described in Section III, remanufacturing and recycling operations can be effectively designed. Through integration with the AAS, PoPan serves as a digital shadow for the EV battery, accompanying it throughout the entire lifecycle. Such an integration enhances data management and data sharing between involved stakeholders.
From the PoPAN presented in Fig. 6, we can query the set of operations to disassemble the battery in the robotic workcells. We can see that the Last Product is the Lid, which has to be manipulated/removed from the EV battery. However, the conditioned edge between the Lid product and Manipulation in this PoPAN requires to evaluate the edges connected to the vertex Manipulation. Since we are currently disassembling, we cannot proceed to the process Screwing. The search algorithm proceeds in the full arrow direction to the Bolts M6 with assigned Screwing process. When finished, it can continue to the Manipulation process, which was previously skipped. Then the search algorithm can proceed to the product Battery Box, which has more than one incoming edges. Prior executing the assigned process Manipulation, the incoming edges have to be passed. Therefore, the search algorithm proceeds to the remaining branches and starts from the vertex that has all dependencies satisfied. During this systematic search, it goes through the entire graph until all processes are done to get to the Initial Product, which is the empty battery box in this case.
Through the performed evaluation, we have confirmed efficacy of PoPAN, as it seamlessly integrates all tasks developed in this use-case. Based on this lessons-learned use-case, we believe PoPAN can be also well-suited for other use-cases.
@@ VI Conclusion and Future Work
This paper proposes a specific adaptation of the PPR Asset Network (PAN) paradigm called Product-oriented PPR Asset Network (PoPAN), having the product structure as the core basis of the descriptive model. Such an adaptation enables to express not only the assembling processes, but also the disassembling ones in just one compact form. The PoPAN model can be afterwards queried to get the right sequence of production processes/operations for various production, reparation, remanufacturing, or recycling purposes during later phases of the whole product life-cycle.
The PoPAN model solves as a model-driven engineering backbone for process management within the whole product life-cycle. It enables large-scale system integration and scalability of the models and respective systems. The proposed approach is demonstrated on a case study on disassembling EV batteries to upcycle them from the automotive applications towards stationary battery storages in industry or households.
Addressing the research question RQ1, this paper proposes a PPR-based model coming up from the PAN model, which however primary reflects the product structure. The model allows to seamlessly combine the production and remanufacturing/disassembling processes into one unified description, addressing the research question RQ2. This overall model was presented in Sec. III.
During the presented research, we were aware of the necessity to foster adaptability of this new modeling approach by industrial and academic stakeholders. Therefore, we have from the very beginning considered export to the AutomationML data format and Asset Administration Shell, addressing the research question RQ3, providing the serialization of the model into AutomationML in Sec. IV.
In the future work, we would like to implement a user-friendly search algorithm, which would generate the right directed acyclic graph of processes for the queried situation and export it into BPMN.

### v2

@@ IV Representation of PoPAN in AutomationML
The AutomationML representation aims at capturing the PoPAN network as a graph. First, we specified the InterfaceClassLib with the Interface Class EdgeVertexInterface. This interface allows the connection of edges and vertices via AutomationML Internal Links. The Interface Class has an attribute Direction, which can have values In for the incoming edge, Out for the outgoing edge, and InOut for the undirected edge. Since PoPAN distinguishes assembly and disassembly directions, the EdgeVertexInterface Class has also attribute Arrow, which can have values OpenArrow for the assembly direction, FullArrow for the disassembly direction, or None for none arrow.
@@
In the third step, we have defined exemplary/generic System Unit Class, which are depicted in Fig. 2. The figure shows four Internal Elements (IE) from the System Unit Class, in compliance with their Roles from the Role Class Library. The Internal Elements are as a matter of fact instances of the aforementioned Role Classes, accompanied by EdgeVertexInterface to enable the PoPAN representation.
@@
The IE Product type represents physical products within the whole product. Each IE Product is uniquely identified by an Id and has a Type including ElementaryProduct, SubProduct, FastenerProduct, InitialProduct, or LastProduct. Additionally, IE Product has assigned Roles Product and Submodel (to correspond with AAS). The IE Process type represents processes performed within the assembly/disassembly operations. IE Process is uniquely identified by an Id, has assigned Role Process, and has a Type. The IE Resource type represents resources needed to perform processes, such as tools and stations. Similar to IE Process, each IE Resource has a unique identifier Id, has an assigned Role Process, and has a Type. The IE Edge type represents connections between PPR components within the whole product. IE Edge is uniquely identified by Id, the Role Edge and has a Type that can have values Physical to represent the physical continuity of products, Structural for the structural continuity of products, LinkingP2P for linking a product to a process, SequenceProcess to represent dependencies/sequences of processes, and LinkingP2R for linking a process to a resource. For all four IE Product types, the EdgeVertexInterface is defined to ensure connection to other IEs within the PoPAN.
@@
The IE Product supports the Role Submodel from the Asset Administration Shell Role Class Library [^3] Online: https://www.automationml.org/wp-content/uploads/2022/04/Asset-Administration-Shell-Representation-V1_0_0.zip. In this place we should note that we are missing a “SubModelElement/Entity” in this AutomationML library, which would be useful for transforming the description into the AASX format.
@@
The example of a generic product structure according to the PoPAN model from the previous section is depicted in Fig. 3. If we need to assign more than one edge to a vertice, we have to increase the number of EdgeVertexInterfaces. In Fig. 4, a network of PoPAN elements represented in AutomationML is shown for a Product 1 from a generic product (Sec. III).
@@
AutomationML supports annotating objects relevant to the Asset Administration Shell (AAS). AAS is becoming an important industrial artifact to be exchanged in the entire supply chain. For creating the AAS, we are using the aforementioned tool AASX Package Explorer, which can import an AutomationML file and transform it into the AASX format, which can accompany the product during its entire life-cycle.
@@
This comprehensive product description, when used in conjunction with the AAS, can accompany the product throughout its entire life-cycle. It can assist in scenarios where defective parts need replacement, while also facilitating the tracking of new or missing products (e.g., a missing screw) to streamline the recycling process.
@@ V Electric Vehicle Battery Remanufacturing Use-Case
In our laboratory conditions, utilizing the Industry 4.0 Testbed [^4] Online: https://www.ciirc.cvut.cz/teams-labs/testbed/ at CTU in Prague – CIIRC, we employ KUKA robots to manipulate the real EV battery according to predefined parameters (see Fig. 5). The primary goal of this use-case is to disassemble the entire EV battery with a flexible robotic system. This approach offers the potential use of modules from the EV battery for secondary applications. Additionally, the secondary goal involves replacing broken modules, which necessitates a combination of assembling and disassembling operations. We were seeking a formalism that allows to accommodate all possible operations in our use-case, and PoPAN fits our requirements perfectly for all considered applications. The visualization of PoPAN for a simplified description of the EV battery is depicted in Fig. 6. It illustrates all relevant data for creating assembly/disassembly operations. As outlined in Section III, this visualization contains products, processes, and resources interconnected by different types of edges, with two types of arrows indicating assembly/disassembly orientations. Following the rules described in Section III, remanufacturing and recycling operations can be effectively designed. Through integration with the AAS, PoPAN serves as a digital shadow for the EV battery, accompanying it throughout the entire life-cycle. Such an integration enhances data management and data sharing between involved stakeholders.
@@
From the PoPAN presented in Fig. 6, we can query the set of operations to disassemble the battery in the robotic workcells. We can see that the Last Product is the Lid, which has to be manipulated/removed from the EV battery. However, the conditioned edge between the Lid product and Manipulation in this PoPAN requires to evaluate the edges connected to the vertex Manipulation. Since we are currently disassembling, we cannot proceed to the process Screwing. The search algorithm proceeds in the full arrow direction to the Bolts M6 with the assigned Screwing process. When finished, it can continue to the Manipulation process, which was previously skipped. Then the search algorithm can proceed to the product Battery Box, which has more than one incoming edge. Prior to executing the assigned process Manipulation, the incoming edges have to be passed. Therefore, the search algorithm proceeds to the remaining branches and starts from the vertex that has all dependencies satisfied. During this systematic search, it goes through the entire graph until all processes are done to get to the Initial Product, which is the empty battery box in this case.
@@
Through the performed evaluation, we have confirmed the efficacy of PoPAN, as it seamlessly integrates all tasks developed in this use-case. Based on this lessons-learned use-case, we believe PoPAN can be also well-suited for other use-cases.
@@ VI Conclusion and Future Work
This paper proposes a specific adaptation of the PAN paradigm called Product-oriented PPR Asset Network (PoPAN), having the product structure as the core basis of the descriptive model. Such an adaptation enables to express not only the assembling processes but also the disassembling ones in just one compact form. The PoPAN model can be afterward queried to get the right sequence of production processes/operations for various production, reparation, remanufacturing, or recycling purposes during later phases of the whole product life-cycle.
The PoPAN model serves as a model-driven engineering backbone for process management within the whole product life-cycle. It enables large-scale system integration and scalability of the models and respective systems. The proposed approach is demonstrated in a case study on disassembling EV batteries to upcycle them from automotive applications towards stationary battery storages in industry or households.
Addressing the research question RQ1, this paper proposes a PPR-based model coming up from the PAN model, which however primarily reflects the product structure. The model allows to seamlessly combine the production and remanufacturing/disassembling processes into one unified description, addressing the research question RQ2. This overall model was presented in Sec. III.
During the presented research, we were aware of the necessity to foster the adaptability of this new modeling approach by industrial and academic stakeholders. Therefore, we have from the very beginning considered an export to the AutomationML data format and Asset Administration Shell, addressing the research question RQ3, providing the serialization of the model into AutomationML in Sec. IV.
In future work, we would like to implement a user-friendly search algorithm for generating the right directed acyclic graph of processes for the queried situation and export it into BPMN.